# High-Dimensional Independence Testing via Maximum and Average Distance Correlations

### v2

@@ Abstract
This paper introduces and investigates the utilization of maximum and average distance correlations for multivariate independence testing. We characterize their consistency properties in high-dimensional settings with respect to the number of marginally dependent dimensions, assess the advantages of each test statistic, examine their respective null distributions, and present a fast chi-square-based testing procedure. The resulting tests are non-parametric and applicable to both Euclidean distance and the Gaussian kernel as the underlying metric. To better understand the practical use cases of the proposed tests, we evaluate the empirical performance of the maximum distance correlation, average distance correlation, and the original distance correlation across various multivariate dependence scenarios, as well as conduct a real data experiment to test the presence of various cancer types and peptide levels in human plasma.
Keywords: unbiased distance correlation, chi-square test, testing independence
@@ 1 Introduction
Given pairs of observations $(x_{i},y_{i})\in\mathbb{R}^{p}\times\mathbb{R}^{q}$ for $i=1,\ldots,n$, assume they are independently identically distributed as $F_{XY}$. The statistical hypothesis for testing independence is formulated as:
@@
Traditional correlation measures like Pearson’s correlation (Pearson,, 1895) are commonly used but unable to detect nonlinear and high-dimensional dependencies. Recent measures, such as the distance correlation (Szekely et al.,, 2007; Szekely and Rizzo,, 2009) and the Hilbert-Schmidt independence criterion (Gretton et al.,, 2005; Gretton and Gyorfi,, 2010), can uncover any type of dependency given sufficient sample size, are zero if and only if independence, share similar characteristics, and can reliably test independence for any joint distribution of any fixed dimensionality. Dependence measures are valuable in various statistical applications, such as feature screening (Li et al.,, 2012; Zhong and Zhu,, 2015; Shen et al.,, 2024), time-series (Zhou,, 2012; Fokianos and Pitsillou,, 2018; Shen et al.,, 2023), conditional independence (Fukumizu et al.,, 2007; Szekely and Rizzo,, 2014; Wang et al.,, 2015), clustering (Szekely and Rizzo,, 2005; Rizzo and Szekely,, 2010), graph testing (Lee et al.,, 2019; Xiong et al.,, 2022), and deep learning (D. Guo and Zha,, 2022; Zhen et al.,, 2022).
Detecting multivariate dependencies, and especially in high-dimensional scenarios, remains a challenging task with limited understanding. As the number of dimensions increases relative to the sample size, the testing power of existing dependence measures may diminish (Shen et al.,, 2020; Ramdas et al.,, 2015). In situations where the dimensions ($p$ or $q$) approach infinity and the sample size grows more slowly than the dimension, distance correlation may fail to detect certain multivariate dependencies (Zhu et al.,, 2019). To address this issue, several solutions have been proposed, such as computing marginal distance covariance for each dimension in $X$ and $Y$ and averaging them to form the test statistic (Zhu et al.,, 2019), or considering a random rotation of the average distance covariance (Huang and Huo,, 2017). Moreover, hypothesis testing through marginal covariance typically involves employing a standard permutation test to calculate the p-value, which can be computationally intensive and cost as much as $\mathcal{O}(rpqn\log(n))$, where $r$ is the number of random permutations.
In this paper, we examine the utilization of maximum distance correlation and average distance correlation as test statistics for multivariate dependence testing. We formulate the maximum distance correlation and the average distance correlation based on pairwise, unbiased, and marginal distance correlations. To understand their respective advantages, we establish their consistency properties for high-dimensional independence testing, using the concept of marginally dependent dimensions. Subsequently, we analyze their limiting null distribution and propose a valid chi-square-based test to calculate p-values, which is significantly faster than permutation tests for large datasets. Our numerical study compares the performance of maximum, average, and original distance correlation using both Euclidean distance and the Gaussian kernel across various simulation settings. Finally, we provide a real data experiment on cancer types and peptide levels in human plasma to illustrate their practical applications. All theorem proofs are in the Appendix.
@@ 2 Background
In this section, we provide a review of existing results, including the unbiased distance correlation, its relationship with the Hilbert-Schmidt independence criterion (HSIC), its validity and consistency in testing independence, the permutation test, the limiting null distribution, and the chi-square test.
We denote the paired sample data, assumed to be independently and identically distributed as $F_{XY}$, as follows:
@@
Moreover, we always assume finite moments for $F_{XY}$ throughout this paper.
Given a distance metric $d(\cdot,\cdot)$ such as the Euclidean metric, we denote $\mathbf{D}^{\mathbf{X}}$ as the $n\times n$ pairwise distance matrix of $\mathbf{X}$ with $\mathbf{D}^{\mathbf{X}}_{ij}=d(x_{i},x_{j})$. Similarly, we denote $\mathbf{D}^{\mathbf{Y}}$ as the pairwise distance matrix of $\mathbf{Y}$.
Next, we compute a modified matrix $\mathbf{C}^{\mathbf{X}}$ as follows:
@@
${\sc\texttt{Dcov}}_{n}(\mathbf{X},\mathbf{Y})=\frac{1}{n(n-3)}{\operatorname{trace}\!\left(\mathbf{C}^{\mathbf{X}}\mathbf{C}^{\mathbf{Y}}\right)},$
${\sc\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})=\frac{{\sc\texttt{Dcov}}_{n}(\mathbf{X},\mathbf{Y})}{\sqrt{{\sc\texttt{Dcov}}_{n}(\mathbf{X},\mathbf{X}){\sc\texttt{Dcov}}_{n}(\mathbf{Y},\mathbf{Y})}}\in[-1,1].$
@@
The above unbiased statistic was introduced in Szekely and Rizzo, (2014). Comparing to the biased statistic introduced in Szekely et al., (2007), the unbiased statistic satisfies
$\mathbb{E}({\sc\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y}))=0$
@@
By default, distance correlation utilizes the Euclidean distance as its metric. However, it is versatile enough to accommodate any distance metric or kernel choice by setting $\mathbf{D}^{\mathbf{X}}$ and $\mathbf{D}^{\mathbf{Y}}$ as two sample kernel matrices. It is worth noting that when the Gaussian kernel is used, distance correlation effectively becomes equivalent to HSIC. In fact, one can interchange between distance and kernel metrics through an appropriate kernel-to-distance transformation (Sejdinovic et al.,, 2013; Shen and Vogelstein,, 2021).
@@
When the metric used is of strong negative type (Lyons,, 2013, 2018), such as the Euclidean distance, or when a characteristic kernel is used (Gretton et al.,, 2005; Fukumizu et al.,, 2007; Gretton and Gyorfi,, 2010), like the Gaussian kernel, the resulting distance correlation exhibits the following property:
${\sc\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}0$
@@
if and only if $X$ and $Y$ are independent. Equivalently, distance correlation converges to a positive constant if and only if $X$ and $Y$ are dependent. This fundamental property makes distance correlation a valid and universally consistent statistic for testing independence using the permutation test, which is a standard approach in testing independence (Good,, 2005; Heller et al.,, 2013; Shen et al.,, 2020).
@@
Specifically, when $X$ and $Y$ are dependent, performing a permutation test on distance correlation yields an asymptotic p-value of 00, leading to testing power that converges to $1$ as the sample size $n$ increases. Conversely, when $X$ and $Y$ are independent, the p-value follows a uniform distribution in the range $[0,1]$, and the testing power equals the type I error level $\alpha$.
The permutation test can be computationally intensive, requiring the permutation of sample data for at least $r=100$ times, with the computation of the permuted statistic for each permutation. Given that distance correlation is typically $O(n^{2})$, the testing procedure entails a computational complexity of $O(rn^{2})$.
Recent advancements have led to a faster implementation of distance correlation with a time complexity of $\mathcal{O}(n\log(n))$ when $p=q=1$ and Euclidean distance is used (Huo and Szekely,, 2016; Chaudhuri and Hu,, 2019). Furthermore, significant progress has been made in characterizing the null distribution of distance correlation, i.e., the distribution when $X$ and $Y$ are independent. This distribution can be fully specified in the limit (Zhang et al.,, 2018) and approximated by the following chi-square distribution (Shen et al.,, 2022):
Theorem 1.
The limiting null distribution of the unbiased distance correlation satisfies
@@
$n\cdot{\sc\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})$ $\stackrel{{\scriptstyle D}}{{\rightarrow}}\sum\limits_{i,j=1}^{\infty}w_{ij}(\mathcal{N}_{ij}^{2}-1),$
@@
where the weights satisfy $w_{ij}\in[0,1]$ and $\sum\limits_{i,j=1}^{\infty}w_{ij}^{2}=1$, and $\mathcal{N}_{ij}$ are independent standard normal distribution.
As a result, for sufficiently large $n$, there exists $\alpha>0$ such that
$\mathcal{N}(0,2)\preceq_{\alpha}n\cdot{\sc\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})\preceq_{\alpha}\chi_{1}^{2}-1$
regardless of the metric choice or marginal distributions.
Definition 1.
$F_{V}(z)\geq F_{U}(z)$
The proof of Theorem 1 can be found in the main theorem of Shen et al., (2022). It holds under the assumption of independence between $X$ and $Y$, finite-moments of $F_{XY}$, an increasing sample size $n$, and regardless of $p,q$ nor the specific distribution of $F_{XY}$.
As a result of Theorem 1, we can use the chi-square test to compute the p-value instead of the permutation test. The p-value calculation is given by:
$\mbox{pval}=1-F_{\chi^{2}_{1}-1}(n\cdot{\sc\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})),$
which has a time complexity of $O(1)$ and is straightforward to implement in any programming language. This test is valid for any type 1 error level $\alpha$ where the upper-tail dominance holds, and it remains universally consistent against any dependence, meaning that the p-value converges to 00 when $X$ and $Y$ are dependent.
Although the level $\alpha$ cannot be exactly determined in closed-form, for sample sizes $n>30$, the chi-square test on distance correlation yields testing power similar to the permutation test and is approximately valid for any $\alpha\leq 0.05$. For example, in high-dimensional scenarios where both $n$ and $p$ tend to infinity, assuming that $X$ is continuous and each dimension of $X$ is exchangeable, the null distribution converges to $\mathcal{N}(0,2)$ (Szekely and Rizzo,, 2013). In such cases, the chi-square test is strictly valid for any $\alpha<0.0875$ (Shen et al.,, 2022).
@@ 3.1 Maximum and Average Distance Correlations
Given $\mathbf{X}=\{x_{i}\in\mathbb{R}^{p}\mbox{ for }i=1,\ldots,n]\}$ as the sample data, let $\mathbf{X}^{s}=\{x_{i}^{s}\in\mathbb{R}\mbox{ for }i=1,\ldots,n\}$ denote the $s$th-dimension of the sample data. Similarly for $\mathbf{Y}$ and $\mathbf{Y}^{t}$ for each $t\in[q]$.
For every $s\in[p]$ and $t\in[q]$, we refer to the distance correlation between $(\mathbf{X}^{s},\mathbf{Y}^{t})$ as the marginal distance correlation, denoted by ${\sc\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})$. To distinguish, we will refer to ${\sc\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})$, which incorporates all dimensions of the sample data, as the original distance correlation.
@@
${\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})=\max\limits_{s\in[p],t\in[q]}{\sc\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t}),$
${\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})=\sum\limits_{s\in[p],t\in[q]}{\sc\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})/pq.$
In essence, ${\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})$ is the maximum of all marginal distance correlations per dimension, while ${\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})$ is the average of all marginal distance correlations. Note that all the marginal sample statistics employ the computation of unbiased distance correlation.
@@ 3.2 Consistency for Testing Marginal Dependence
Clearly, $\delta(X,Y)\in[0,pq]$. If $\delta(X,Y)>0$, it implies that $X$ and $Y$ must be dependent. However, the reverse is not always true; that is, $X$ and $Y$ may be dependent, while $\delta(X,Y)$ could be 00. In practice, creating a counter-example usually requires special construction, and the concept of marginal dependence effectively captures a significant portion of dependence.
@@
${\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})$ $\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}0$
regardless of $pq$.
On the other hand, the maximum distance correlation satisfies:
${\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})$ $\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}0$
when $pq=o(e^{\sqrt{n}})$.
@@
${\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})>0$
regardless of $pq$.
On the other hand, the average distance correlation satisfies:
${\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}c\geq 0,$
@@
The maximum distance correlation is asymptotically consistent in testing the existence of marginal dependence when $pq$ is fixed or $pq=o(e^{\sqrt{n}})$ for increasing $pq$.
Therefore, as long as $n$ is not too small, the maximum correlation is more advantageous because the average statistic may be too small when a dependence signal is present in only a few dimensions. On the other hand, for small-sample problems, the maximum correlation can be significantly biased under the null, potentially inflating its p-value. These behaviors are also observed in the numerical study.
3.3 Limiting Null Distribution and Chi-square-based Tests
While one could employ the permutation test on either the maximum or average statistic, it tends to be very slow for large $n$. To expedite the testing process, we delve into the null distributions of the maximum and average statistics:

### v3

@@ Abstract
This paper investigates the utilization of maximum and average distance correlations for multivariate independence testing. We characterize their consistency properties in high-dimensional settings with respect to the number of marginally dependent dimensions, compare the advantages of each test statistic, examine their respective null distributions, and present a fast chi-square-based testing procedure. The resulting tests are non-parametric and applicable to both Euclidean distance and the Gaussian kernel as the underlying metric. To better understand the practical use cases of the proposed tests, we evaluate the empirical performance of the maximum distance correlation, average distance correlation, and the original distance correlation across various multivariate dependence scenarios, as well as conduct a real data experiment to test the presence of various cancer types and peptide levels in human plasma.
keywords:
unbiased distance correlation, chi-square test, testing independence
@@ 1 Introduction
Given pairs of observations $(x_{i},y_{i})\in\mathbb{R}^{p}\times\mathbb{R}^{q}$ for $i=1,\ldots,n$, assume they are independently identically distributed as $F_{XY}$ of finite moments. The statistical hypothesis for testing independence is formulated as:
@@
Traditional correlation measures like Pearson’s correlation [1] are commonly used but unable to detect nonlinear and high-dimensional dependencies. Recent measures, such as the distance correlation [2, 3] and the Hilbert-Schmidt independence criterion [4, 5], can uncover any type of dependency given sufficient sample size, are asymptotically zero if and only if independence, and can consistently test independence for any joint distribution of fixed dimensionality. Due to such consistency properties, dependence measures have become widely used in various applications, such as feature screening [6, 7, 8, 9], time-series [10, 11, 12], conditional independence [13, 14, 15], K-sample test [16, 17], clustering [18, 19], graph testing [20, 21], and deep learning [22, 23].
Detecting multivariate dependencies, and especially in high-dimensional scenarios, remains a challenging task with limited understanding. As the number of dimensions increases relative to the sample size, the testing power of existing dependence measures may diminish [24]. In situations where the dimensions ($p$ or $q$) approach infinity and the sample size grows more slowly than the dimension, distance correlation may fail to detect certain multivariate dependencies [25]. To address this issue, a common solution involves projecting the high-dimensional data into a low-dimensional space and then considering the marginal dependence, such as the average marginal covariance [25], projection correlation [26], and average covariance of random rotations [27].
In this paper, we examine the utilization of maximum distance correlation and average distance correlation as sample test statistics for multivariate dependence testing. We formulate the maximum distance correlation and the average distance correlation based on pairwise, unbiased, and marginal distance correlations. To understand their respective advantages, we establish their consistency properties for high-dimensional independence testing, using the concept of marginally dependent dimensions. Subsequently, we analyze their limiting null distribution, and propose a valid chi-square-based test for computing p-values, which is significantly faster than the standard permutation test. Our numerical study compares the performance of maximum, average, and original distance correlation using both Euclidean distance and the Gaussian kernel across various simulation settings. Finally, we provide a real data experiment on cancer types and peptide levels in human plasma to illustrate their practical applications. All theorem proofs and additional simulations are in the appendix.
@@ 2 Background
In this section, we provide a review of existing results, including the unbiased distance correlation, its relationship with the Hilbert-Schmidt independence criterion (HSIC), its validity and consistency in testing independence, the standard permutation test, the limiting null distribution, and the chi-square test. We denote the paired sample data as follows:
@@
Namely, $(\mathbf{X},\mathbf{Y})$ represents sample realizations of the random variable $(X,Y)$, where each sample pair $(x_{i},y_{i})$ is assumed to be independently and identically distributed as $F_{XY}$. Throughout this paper, we assume finite moments for $F_{XY}$.
Distance Correlation
Given a distance metric $d(\cdot,\cdot)$ such as the Euclidean metric, we denote $\mathbf{D}^{\mathbf{X}}$ as the $n\times n$ pairwise distance matrix of $\mathbf{X}$ with $\mathbf{D}^{\mathbf{X}}_{ij}=d(x_{i},x_{j})$. Similarly, we denote $\mathbf{D}^{\mathbf{Y}}$ as the pairwise distance matrix of $\mathbf{Y}$. Next, we compute a modified matrix $\mathbf{C}^{\mathbf{X}}$ as follows:
@@
${\texttt{Dcov}}_{n}(\mathbf{X},\mathbf{Y})=\frac{1}{n(n-3)}{\operatorname{trace}\!\left(\mathbf{C}^{\mathbf{X}}\mathbf{C}^{\mathbf{Y}}\right)},$
${\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})=\frac{{\texttt{Dcov}}_{n}(\mathbf{X},\mathbf{Y})}{\sqrt{{\texttt{Dcov}}_{n}(\mathbf{X},\mathbf{X}){\texttt{Dcov}}_{n}(\mathbf{Y},\mathbf{Y})}}\in[-1,1].$
@@
Theoretical Property
The above unbiased statistic was introduced in [14]. Comparing to the biased statistic introduced in [2], the unbiased statistic satisfies
$\mathbb{E}({\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y}))=0$
@@
By default, distance correlation utilizes the Euclidean distance as its metric. Nevertheless, it may accommodate any distance metric or kernel choice by setting $\mathbf{D}^{\mathbf{X}}$ and $\mathbf{D}^{\mathbf{Y}}$ as the corresponding distance or kernel matrices. It is worth noting that when the Gaussian kernel is used, distance correlation effectively becomes equivalent to HSIC. In fact, one can interchange between distance and kernel metrics through an appropriate kernel-to-distance transformation [28, 29].
@@
When the metric used is of strong negative type [30, 31], such as the Euclidean distance, or when a characteristic kernel is used [4, 13], like the Gaussian kernel, the resulting distance correlation exhibits the following property:
${\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}0$
@@
if and only if $X$ and $Y$ are independent. This fundamental property makes distance correlation a valid and universally consistent statistic for testing independence.
Hypothesis Test
To conduct testing on sample data, one needs to approximate the null distribution of the test statistic, typically achieved through a standard permutation test [32]. Given $\mathbf{X}$ and $\mathbf{Y}$, the process involves randomly permuting the indices of $\mathbf{Y}$ for $R$ iterations. The p-value is then computed as:
$\mbox{pval}=\frac{1}{R}\sum_{r=1}^{R}1\{{\texttt{Dcor}}_{n}(\mathbf{X},\pi_{r}(\mathbf{Y}))>{\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})\}$
@@
where each $\pi_{r}$ represents a random permutation of size $n$. Specifically, when $X$ and $Y$ are dependent, performing a permutation test on distance correlation yields an asymptotic p-value of 00, leading to testing power that converges to $1$ as the sample size $n$ increases. Conversely, when $X$ and $Y$ are independent, the p-value follows a Uniform distribution in the range $[0,1]$, and the testing power equals the type I error level $\alpha$. The primary drawback of the permutation test is its computational complexity, which requires repeated computation of the statistic for $R$ iterations, often set to $100$ or more, rendering it time-consuming for large sample sizes.
Recent works have significantly reduced the computational cost of distance correlation [33, 34] and the resulting permutation test [35, 9]. Specifically, the null distribution of distance correlation satisfies the following [36]:
@@
$n\cdot{\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})$ $\stackrel{{\scriptstyle D}}{{\rightarrow}}\sum\limits_{i,j=1}^{\infty}w_{ij}(\mathcal{N}_{ij}^{2}-1),$
@@
where the weights satisfy $w_{ij}\in[0,1]$ and $\sum_{i,j=1}^{\infty}w_{ij}^{2}=1$, and $\mathcal{N}_{ij}$ are independent standard normal distributions. The null distribution can be further bounded by a chi-square-based distribution, enabling the use of the following chi-square test:
$\mbox{pval}=1-F_{\chi^{2}_{1}-1}(n\cdot{\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})).$
This approximation is approximately valid for any $\alpha<0.0875$ [35], and has a time complexity of $O(1)$ and is straightforward to implement in any programming language.
Difference from Mutual Independence
A related but fundamentally distinct concept is that of testing mutual independence within a random vector $X$. For instance, a recent paper [25] introduced a similar idea of using the maximum of marginal statistics to construct a more powerful test for mutual independence. However, testing independence between $X\in\mathbb{R}^{p}$ and $Y\in\mathbb{R}^{q}$ is conceptually and technically different from testing mutual independence among the marginals of $X$. While there are similarities in the construction of test statistics (e.g., via marginal aggregations), methods developed for one problem are generally not applicable to the other.
To be precise, testing mutual independence within $X$ corresponds to the null hypothesis:
$H_{0}:F_{X}=\prod_{j=1}^{p}F_{X^{j}},$
i.e., all coordinates of $X$ are mutually independent. This is a much stronger requirement than pairwise independence, and implies that there is no interaction or shared randomness among any subset of the components of $X$.
By contrast, testing independence between $X$ and $Y$ involves the hypothesis:
$H_{0}:F_{XY}=F_{X}\cdot F_{Y},$
which concerns the relationship between two random vectors, regardless of any dependence structure within $X$ or within $Y$ individually.
To illustrate the distinction, consider the combined vector $Z=[X,Y]$. If $X\perp Y$, this does not imply that the components of $X$ are mutually independent — hence, $Z$ may or may not be mutually independent. Conversely, if $X$ and $Y$ are dependent, then $Z$ must be mutually dependent. As a result, applying a mutual independence test to $Z$ is not a valid or consistent method for testing whether $X\perp Y$, because mutual dependence of $Z$ does not uniquely determine the presence or absence of dependence between $X$ and $Y$. The reverse also holds: tests for independence between $X$ and $Y$ are not valid nor consistent for assessing mutual independence within $X$.
The only special case in which the two problems coincide is when both $X$ and $Y$ are univariate. In that setting, testing independence between $X$ and $Y$ is equivalent to testing mutual independence within $Z=[X,Y]$, which forms the basis for using marginal dependence measures for mutual independence testing in the bivariate case, as in [25]. Outside of this setting, however, the two problems remain fundamentally different.
Therefore, while both settings involve notions of statistical dependence, they target distinct structural properties of the data, and the associated testing methodologies are not interchangeable in general.
@@ 3.1 Maximum and Average Distance Correlations
Given $\mathbf{X}=\{x_{i}\in\mathbb{R}^{p}\mbox{ for }i=1,\ldots,n]\}$ as the sample data, let $\mathbf{X}^{s}=\{x_{i}^{s}\in\mathbb{R}\mbox{ for }i=1,\ldots,n\}$ denote the $s$th-dimension of the sample data. Similarly for $\mathbf{Y}$ and $\mathbf{Y}^{t}$ for each $t\in[q]$. For every $s\in[p]$ and $t\in[q]$, we refer to the distance correlation between $(\mathbf{X}^{s},\mathbf{Y}^{t})$ as the marginal distance correlation, denoted by ${\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})$. To distinguish, we will refer to ${\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})$, which incorporates all dimensions of the sample data, as the original distance correlation.
@@
${\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})=\max\limits_{s\in[p],t\in[q]}{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t}),$
${\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})=\sum\limits_{s\in[p],t\in[q]}{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})/pq.$
In essence, ${\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})$ is the maximum of all marginal distance correlations per dimension, while ${\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})$ is the average of all marginal distance correlations. Note that in this paper, all the sample statistics utilize the unbiased distance correlation.
@@ 3.2 Consistency for Testing Marginal Dependence
Clearly, $\delta(X,Y)\in[0,pq]$. If $\delta(X,Y)>0$, it implies that $X$ and $Y$ must be dependent. However, the reverse is not always true; that is, $X$ and $Y$ may be dependent, while $\delta(X,Y)$ could be 00. One such example is provided in appendix A4. In practice, creating such examples requires special construction, and the concept of marginal dependence captures a significant portion of dependence.
@@
${\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})$ $\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}0$
regardless of $pq$. On the other hand, the maximum distance correlation satisfies:
${\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})$ $\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}0$
when $pq=o(\sqrt{n}e^{n})$.
@@
${\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}c>0$
regardless of $pq$. On the other hand, the average distance correlation satisfies:
${\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}c\geq 0,$
@@
The maximum distance correlation is asymptotically consistent in testing the existence of marginal dependence when $pq$ is fixed or $pq=o(\sqrt{n}e^{n})$ for increasing $pq$.
Therefore, the maximum correlation is advantageous when the dependence signal is sparse, i.e., concentrated in a few marginal pairs. In contrast, the average statistic is more powerful when there are many weak associations spread across many dimensions.
3.3 Limiting Null Distribution
While one could employ the permutation test on either the maximum or average statistic (detailed procedure in appendix), it tends to be very slow for large $n$. To expedite the testing process, we consider the null distributions of the maximum and average statistics: